If the force between two charged objects is to be left unchanged even though the charge on one of the objects is halved keeping the other the same the original distance of sepration d should be equal to

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear pallavi

let the force on charged object is q and Q

F=1/4∏ε  qQ/d

 now let the q become q/2    and seperation now x

F= 1/4∏ε (q/2)Q/x

so 1/4∏ε  qQ/d2   = 1/4∏ε (q/2)Q/x

  1/d2 =1/2x2

  x = d/√2

so  seperation between them should be  decrese by = d-d/√2

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin